I have an equilateral triangle with side = x. The area is 5000 sq metres
Need a formula to work out x.
Thanks

Answer by scott8148(6628)

the altitude of an equilateral triangle divides the triangle into two 30-60-90 right triangles
___ if the side is s, then the height is (s√3)/2

so, in this case ___ A = (1/2)bh ___ 5000 = (1/2)x[(x√3)/2]